# PTQ4DiT: Post-training Quantization for Diffusion Transformers

### v1

@@ Abstract
The recent introduction of Diffusion Transformers (DiTs) has demonstrated exceptional capabilities in image generation by using a different backbone architecture, departing from traditional U-Nets and embracing the scalable nature of transformers. Despite their advanced capabilities, the wide deployment of DiTs, particularly for real-time applications, is currently hampered by considerable computational demands at the inference stage. Post-training Quantization (PTQ) has emerged as a fast and data-efficient solution that can significantly reduce computation and memory footprint by using low-bit weights and activations. However, its applicability to DiTs has not yet been explored and faces non-trivial difficulties due to the unique design of DiTs. In this paper, we propose PTQ4DiT, a specifically designed PTQ method for DiTs. We discover two primary quantization challenges inherent in DiTs, notably the presence of salient channels with extreme magnitudes and the temporal variability in distributions of salient activation over multiple timesteps. To tackle these challenges, we propose Channel-wise Salience Balancing (CSB) and Spearmen’s -guided Salience Calibration (SSC). CSB leverages the complementarity property of channel magnitudes to redistribute the extremes, alleviating quantization errors for both activations and weights. SSC extends this approach by dynamically adjusting the balanced salience to capture the temporal variations in activation. Additionally, to eliminate extra computational costs caused by PTQ4DiT during inference, we design an offline re-parameterization strategy for DiTs. Experiments demonstrate that our PTQ4DiT successfully quantizes DiTs to 8-bit precision (W8A8) while preserving comparable generation ability and further enables effective quantization to 4-bit weight precision (W4A8) for the first time.
@@ 1 Introduction
Model quantization [30, 29, 25] is a prominent technique for accelerating deep learning models because of its high compression rate and significant reduction in inference time. This technique transforms model weights and activations into low-bit formats, which directly reduces the computational burden and memory usage. Among various methods, Post-training Quantization (PTQ) stands out as a leading approach since it circumvents the need to re-train the original model [57, 41, 18]. Practically, PTQ requires only a small dataset for fast calibration, thus is highly suitable for quantizing DiTs, whose re-training process involves extensive data and computational resources [14, 6].
@@
Recognizing the variability in activations over different timesteps, we further extend the concept of channel salience along the temporal dimension and propose Spearmen’s -guided Salience Calibration (SSC). This method refines the Salience Balancing Matrices to comprehensively evaluate activation salience over timesteps, with more emphasis on timesteps where the complementarity between salient activation and weight channels is more significant. Furthermore, we design a re-parameterization scheme that can offline absorb these Salience Balancing Matrices into adjacent layers, thus avoiding additional computation overhead at the inference stage.
@@ 2.2 Model Quantization
Model quantization is a compression technique that improves the inference efficiency of deep learning models by transforming full-precision tensors into -bit integer approximations, leading to direct computational acceleration and memory saving [30, 57, 8, 25, 19]. Formally, the quantization process can be defined as:
@@
PTQ has been effectively applied to a wide range of neural networks, including CNNs [20, 48, 23], Language Transformers [8, 52, 22], Vision Transformers [57, 13, 21, 26], and U-Net-based Diffusion models [41, 18, 15, 47]. Despite its demonstrated success, PTQ’s applicability to Diffusion Transformers (DiTs) remains unexplored, presenting a significant open challenge within the research community. To bridge this gap, our work delves into the unique challenges of quantizing DiTs and introduces the first PTQ method for DiTs that can fruitfully preserve their generation performance.
@@ 4 PTQ4DiT
To overcome the identified challenges, we propose Channel-wise Salience Balancing (CSB) and Spearmen’s -guided Salience Calibration (SSC) in our PTQ4DiT in Sections 4.1 and 4.2, respectively. Subsequently, we devise a re-parameterization scheme in Section 4.3, eliminating extra computational demands of PTQ4DiT during inference while maintaining the mathematical equivalence.
@@ 4.1 Channel-wise Salience Balancing
4.2 Spearmen’s rho-guided Salience Calibration
@@
To accurately gauge the activation channel salience under multi-timestep scenarios, we propose Spearmen’s -guided Salience Calibration (SSC). This offers a comprehensive evaluation of activation salience, with enhanced focus allocated to the timesteps where the complementarity property is more significant, facilitating effective salience balancing between activation and weight channels. Essentially, the lower the correlation between activation salience and weight salience , the greater reduction effect in overall channel salience (Eq. (8)). The intuition of SSC is visualized in Figure 2 (Right). Mathematically, we formulate the Spearmen’s -calibrated Temporal Salience by selectively aggregating the activation salience along timesteps:
where weighting factors are derived from a normalized exponential form of inverse Spearman’s statistic [44, 51]:
Here, computes the correlation between two sequences, and serves as the weighting factor for activation salience at timestep . In this method, inversely reflects the correlation coefficient , thereby prioritizing timesteps where there is a higher degree of complementarity in salience between activations and weights. Subsequently, we utilize for activation salience in Eqs. (5), (6), and (7), yielding refined Salience Balancing Matrices, denoted as and . By applying SSC, we calibrate the activation salience within CSB to strategically account for the temporal variations during the denoising process. Appendix presents the full Algorithm for PTQ4DiT.
@@ 4.3 Re-Parameterization
The proof is provided in Appendix. Furthermore, we design a re-parameterization scheme for DiTs, allowing for obtaining and without extra computational burden during inference. Specifically, we update the weight matrix of linear layer to offline and seamlessly integrate into the preceding linear transformation operations. This integration includes adaptations to adaLN [35, 34] and matrix multiplications within attention mechanisms [46]. Appendix discusses these adaptations.
@@
Equivalently, we fuse into the MLPs responsible for regressing these parameters, thus avoiding additional computation overhead at inference time. Detailed derivations are provided in Appendix.
@@ 5.2 Quantization Performance
FP models, whereas most baseline methods experience significant performance losses. At the more stringent 4-bit weight precision (W4A8), all baseline methods exhibit more considerable degradation. For instance, under 250 timesteps, PTQ4DM [41] sees a drastic FID increase of 68.05. In contrast, our PTQ4DiT only incurs a slight increase of 2.56. This resilience remains evident as the number of timesteps decreases, underscoring the robustness of PTQ4DiT in resource-limited environments. Moreover, PTQ4DiT markedly outperforms mainstream methods at the higher 512512 resolution, further validating its superiority. For example, using 250 timesteps, PTQ4DiT substantially lowers FID by 41.26 and sFID by 9.83 over the second-best method, Q-Diffusion. Figure 6 depicts the efficiency-vs-efficacy trade-off on W8A8 across various timestep configurations. Our PTQ4DiT achieves comparable performance levels to FP models but with considerably reduced computational costs, offering a viable alternative for high-quality image generation. Figure 5 and Appendix also present randomly generated images for visual comparisons, highlighting PTQ4DiT’s ability to produce images of superior quality.
@@ 6 Conclusion
This paper proposes PTQ4DiT, a novel Post-training Quantization (PTQ) method for Diffusion Transformers (DiTs). Our analysis identifies the primary challenges in effective DiT quantization: the pronounced quantization errors incurred by salient channels with extreme magnitudes and the temporal variability in salient activation. To address these challenges, we design Channel-wise Salience Balancing (CSB) and Spearmen’s -guided Salience Calibration (SSC). Specifically, CSB utilizes the complementarity nature of salient channels to redistribute the extremes within activations and weights toward the balanced salience. SSC dynamically adjusts salience evaluations across different timesteps, prioritizing timesteps where salient activation and weight channels exhibit significant complementarity, thereby mitigating overall quantization difficulties. To avoid extra computational costs of PTQ4DiT, we also devise a re-parameterization strategy for efficient inference. Experiments show that our PTQ4DiT can effectively quantize DiTs to 8-bit precision (W8A8) and further advance to 4-bit weight (W4A8) while maintaining high-quality image generation capabilities.

### v3

@@ Abstract
The recent introduction of Diffusion Transformers (DiTs) has demonstrated exceptional capabilities in image generation by using a different backbone architecture, departing from traditional U-Nets and embracing the scalable nature of transformers. Despite their advanced capabilities, the wide deployment of DiTs, particularly for real-time applications, is currently hampered by considerable computational demands at the inference stage. Post-training Quantization (PTQ) has emerged as a fast and data-efficient solution that can significantly reduce computation and memory footprint by using low-bit weights and activations. However, its applicability to DiTs has not yet been explored and faces non-trivial difficulties due to the unique design of DiTs. In this paper, we propose PTQ4DiT, a specifically designed PTQ method for DiTs. We discover two primary quantization challenges inherent in DiTs, notably the presence of salient channels with extreme magnitudes and the temporal variability in distributions of salient activation over multiple timesteps. To tackle these challenges, we propose Channel-wise Salience Balancing (CSB) and Spearman’s -guided Salience Calibration (SSC). CSB leverages the complementarity property of channel magnitudes to redistribute the extremes, alleviating quantization errors for both activations and weights. SSC extends this approach by dynamically adjusting the balanced salience to capture the temporal variations in activation. Additionally, to eliminate extra computational costs caused by PTQ4DiT during inference, we design an offline re-parameterization strategy for DiTs. Experiments demonstrate that our PTQ4DiT successfully quantizes DiTs to 8-bit precision (W8A8) while preserving comparable generation ability and further enables effective quantization to 4-bit weight precision (W4A8) for the first time.
@@ 1 Introduction
Model quantization [33, 32, 28] is a prominent technique for accelerating deep learning models because of its high compression rate and significant reduction in inference time. This technique transforms model weights and activations into low-bit formats, which directly reduces the computational burden and memory usage. Among various methods, Post-training Quantization (PTQ) stands out as a leading approach since it circumvents the need to re-train the original model [62, 44, 18, 63, 22]. Practically, PTQ requires only a small dataset for fast calibration, thus is highly suitable for quantizing DiTs, whose re-training process involves extensive data and computational resources [14, 6].
@@
Recognizing the variability in activations over different timesteps, we further extend the concept of channel salience along the temporal dimension and propose Spearman’s -guided Salience Calibration (SSC). This method refines the Salience Balancing Matrices to comprehensively evaluate activation salience over timesteps, with more emphasis on timesteps where the complementarity between salient activation and weight channels is more significant. Furthermore, we design a re-parameterization scheme that can offline absorb these Salience Balancing Matrices into adjacent layers, thus avoiding additional computation overhead at the inference stage.
@@ 2.2 Model Quantization
Model quantization is a compression technique that improves the inference efficiency of deep learning models by transforming full-precision tensors into -bit integer approximations, leading to direct computational acceleration and memory saving [33, 62, 8, 28, 19, 64]. Formally, the quantization process can be defined as:
@@
PTQ has been effectively applied to a wide range of neural networks, including CNNs [20, 52, 25], Language Transformers [8, 57, 24, 23, 27], Vision Transformers [62, 13, 21, 29], and U-Net-based Diffusion models [44, 18, 51, 50]. Despite its demonstrated success, PTQ’s applicability to Diffusion Transformers (DiTs) remains unexplored, presenting a significant open challenge within the research community. To bridge this gap, our work delves into the unique challenges of quantizing DiTs and introduces the first PTQ method for DiTs that can fruitfully preserve their generation performance.
@@ 4 PTQ4DiT
To overcome the identified challenges, we propose Channel-wise Salience Balancing (CSB) and Spearman’s -guided Salience Calibration (SSC) in our PTQ4DiT in Sections 4.1 and 4.2, respectively. Subsequently, we devise a re-parameterization scheme in Section 4.3, eliminating extra computational demands of PTQ4DiT during inference while maintaining the mathematical equivalence.
@@ 4.1 Channel-wise Salience Balancing
4.2 Spearman’s -guided Salience Calibration
@@
To accurately gauge the activation channel salience under multi-timestep scenarios, we propose Spearman’s -guided Salience Calibration (SSC). This offers a comprehensive evaluation of activation salience, with enhanced focus allocated to the timesteps where the complementarity property is more significant, facilitating effective salience balancing between activation and weight channels. Essentially, the lower the correlation between activation salience and weight salience , the greater reduction effect in overall channel salience (Eq. (8)). The intuition of SSC is visualized in Figure 2 (Right). Mathematically, we formulate the Spearman’s -calibrated Temporal Salience by selectively aggregating the activation salience along timesteps:
where weighting factors are derived from a normalized exponential form of inverse Spearman’s statistic [47, 55, 56]:
Here, computes the correlation between two sequences, and serves as the weighting factor for activation salience at timestep . In this method, inversely reflects the correlation coefficient , thereby prioritizing timesteps where there is a higher degree of complementarity in salience between activations and weights. Subsequently, we utilize for activation salience in Eqs. (5), (6), and (7), yielding refined Salience Balancing Matrices, denoted as and . By applying SSC, we calibrate the activation salience within CSB to strategically account for the temporal variations during the denoising process. Appendix B presents the full Algorithm for PTQ4DiT.
@@ 4.3 Re-Parameterization
The proof is provided in Appendix C. Furthermore, we design a re-parameterization scheme for DiTs, allowing for obtaining and without extra computational burden during inference. Specifically, we update the weight matrix of linear layer to offline and seamlessly integrate into the preceding linear transformation operations. This integration includes adaptations to adaLN [38, 37] and matrix multiplications within attention mechanisms [49]. Appendix A discusses these adaptations.
@@
Equivalently, we fuse into the MLPs responsible for regressing these parameters, thus avoiding additional computation overhead at inference time. Detailed derivations are provided in Appendix D.
@@ 5.2 Quantization Performance
FP models, whereas most baseline methods experience significant performance losses. At the more stringent 4-bit weight precision (W4A8), all baseline methods exhibit more considerable degradation. For instance, under 250 timesteps, PTQ4DM [44] sees a drastic FID increase of 68.05. In contrast, our PTQ4DiT only incurs a slight increase of 2.56. This resilience remains evident as the number of timesteps decreases, underscoring the robustness of PTQ4DiT in resource-limited environments. Moreover, PTQ4DiT markedly outperforms mainstream methods at the higher 512512 resolution, further validating its superiority. For example, using 250 timesteps, PTQ4DiT substantially lowers FID by 41.26 and sFID by 9.83 over the second-best method, Q-Diffusion. Figure 6 depicts the efficiency-vs-efficacy trade-off on W8A8 across various timestep configurations. Our PTQ4DiT achieves comparable performance levels to FP models but with considerably reduced computational costs, offering a viable alternative for high-quality image generation. Figures 5, 8, and 9 also present randomly generated images for visual comparisons, highlighting PTQ4DiT’s ability to produce images of superior quality.
@@ 6 Conclusion
This paper proposes PTQ4DiT, a novel Post-training Quantization (PTQ) method for Diffusion Transformers (DiTs). Our analysis identifies the primary challenges in effective DiT quantization: the pronounced quantization errors incurred by salient channels with extreme magnitudes and the temporal variability in salient activation. To address these challenges, we design Channel-wise Salience Balancing (CSB) and Spearman’s -guided Salience Calibration (SSC). Specifically, CSB utilizes the complementarity nature of salient channels to redistribute the extremes within activations and weights toward the balanced salience. SSC dynamically adjusts salience evaluations across different timesteps, prioritizing timesteps where salient activation and weight channels exhibit significant complementarity, thereby mitigating overall quantization difficulties. To avoid extra computational costs of PTQ4DiT, we also devise a re-parameterization strategy for efficient inference. Experiments show that our PTQ4DiT can effectively quantize DiTs to 8-bit precision (W8A8) and further advance to 4-bit weight (W4A8) while maintaining high-quality image generation capabilities.
Acknowledgements. This research is supported by NSF IIS-2309073 and ECCS-2123521. This article solely reflects the opinions and conclusions of authors and not funding agencies.
Appendix A Structures of MHSA and PF with Adjusted Linear Layers
The Multi-Head Self-Attention (MHSA) and Pointwise Feedforward (PF) modules are essential for processing input tokens and conditional information in DiT Blocks [37]. As depicted in Figure 7, we incorporate our Channel-wise Salience Balancing (CSB) and Spearman’s -guided Salience Calibration (SSC) techniques into the linear layers within both modules. These techniques are designed to mitigate the quantization difficulties by dynamically adjusting the salience of activations and weights via Salience Balancing Matrices. Through the adjustments, CSB and SSC allow for more uniform distributions of activation and weight magnitudes across salient channels, reducing the impact of extreme values and enhancing the overall stability of the quantization process.
To eliminate additional computational demands during inference, the Salience Balancing Matrices, and , are pre-integrated into the DiT Blocks. Specifically, we replace the weight matrix of the linear layer with and integrate into the preceding linear transformations. For Projection1 and FC1, is absorbed into the MLPs before the adaLN modules, while for Projection2, it can be absorbed within the matrix multiplication [12, 53, 43]. Derivations of the integration are provided in Appendix D.
Appendix B PTQ4DiT Pipeline
This section provides a comprehensive description of the PTQ4DiT pipeline, detailed in Algorithm 1. The PTQ4DiT is designed to enhance the performance of quantized DiTs by addressing quantization challenges through sophisticated salience estimation and balancing strategies.
The full algorithm mainly consists of five steps: ❶ The pipeline begins with estimating activation and weight salience for the pre-trained model using a calibration dataset. ❷ Following the estimation, we employ Spearman’s -guided Salience Calibration to compute correlation coefficients between activation salience and weight salience, which helps determine the weighting factors for each timestep. These factors are crucial for computing a temporally adjusted salience, which aims to minimize quantization errors that typically occur due to misalignment in salience peaks across the timesteps. ❸ The Channel-wise Salience Balancing step follows, wherein Salience Balancing Matrices are constructed to redistribute the activation and weight values channel-wise. Specifically, for each channel , the balanced salience is given by:
where is the -th element of . Then, we formulate the refined Salience Balancing Matrices and based on and , as detailed in Eq. (15). This step is pivotal in aligning the activation and weight distributions, thereby minimizing the overall quantization difficulty. ❹ In the Re-Parameterization phase, these balancing matrices are integrated into the pre-trained model, ensuring that no additional computational cost is required during inference. This integration maintains computational efficiency while retaining the benefits of our salience balancing technique. ❺ Finally, we perform quantization on the model with balanced activations and weights, setting the stage for the deployment of efficient and effective quantized DiTs in resource-constrained environments.
Appendix C Proof of Mathematical Equivalence
In this section, we provide detailed proof demonstrating that our PTQ4DiT maintains mathematical equivalence to the original linear layers. This proof ensures that the balancing operation does not alter the original computational outcomes of the full-precision models.
In PTQ4DiT, we introduce the Salience Balancing Matrices and , which are diagonal matrices intended to balance the salience across activation and weight channels. We verify the inverse relationship of and mentioned in Section 4.3:
where denotes the identity matrix. Therefore, we can derive the mathematical equivalence:
Appendix D Derivations of Post-adaLN Integration
This section details the integration of the Salience Balancing Matrix into the MLPs before the adaptive Layer Norm (adaLN) modules [37], aimed at eliminating extra computational overhead at the inference stage. Recall that the initial formulation of adaLN on the input latent noise is given by:
where are scale and shift parameters, respectively, regressed by MLPs based on the conditional input :
Here, are weight matrices, and are bias terms. In PTQ4DiT, is obtained by applying to the output of adaLN as follows:
which echos with Eq. (13). To avoid additional matrix multiplications in and , we can pre-absorb in MLPs’ weights and biases offline, expressed as:
Thus, the re-parameterized MLPs can directly produce the adjusted scale and shift parameters:
This allows for obtaining without extra computational burden at the inference stage.
Appendix E Additional Visualization Results
Figures 8 and 9 supplement visualization results of our PTQ4DiT on W8A8 quantization, compared with baseline PTQ methods and the full-precision (FP) counterpart, on ImageNet 512512 and 256256. Our method generates results that closely mirror those of the FP models, presenting finer details and richer semantic content than the baseline approaches.
Appendix F Limitations and Broader Impacts
This work introduces a pioneering solution facilitating the broad deployment of Diffusion Transformers (DiTs) through Post-training Quantization. Our method substantially lowers computational and memory demands, thereby improving the accessibility of DiTs. Currently, our research concentrates on visual generation. For future work, we plan to extend our methodology to other generative models across various modalities, such as audio and 3D. However, there remains an inherent risk that these generative models could be utilized to produce disinformation. While our study contributes to the widespread application of DiTs, it does not address such ethical risks. We recognize the importance of developing safeguards and encourage further research into strategies that can prevent the misuse of these powerful generative technologies.